OpenStudy (anonymous):

Eliminate the “z” variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1

OpenStudy (chaise):

2x – y + 2z = –15 2x – 2y – z = 1 <-- multiply every term by two. 2x-y+2z=-15 4x-4y-2z=2. Add the two equations together. 6x-5y+0z=-13 Seeing as though there is 0z it can be disregarded in the equation You are left with 6x-5y=-13.

OpenStudy (anonymous):

Thank you.